# Lecture Notes on Spectral Independence and Bases of a Matroid: Local-to-Global and Trickle-Down from a Markov Chain Perspective

### v3

@@ 1.4 Spectral Independence
1.5 Main Results
Spectral independence was introduced by Anari, Liu, and Oveis Gharan [ALO20]. They proved the following result that spectral independence implies polynomial mixing time of the Glauber dynamics.
Theorem 1 ([ALO20]).
Suppose there exists a constant $\eta>0$ such that the system is $\eta$-spectrally independent, then for any $n$-vertex graph $G=(V,E)$, the relaxation time of the Glauber dynamics satisfies:
[…]
Therefore, the mixing time satisfies $T_{\mathrm{mix}}=O(n^{2+\eta}\log{n})$.
The above result was improved by Chen, Liu, and Vigoda [CLV21] to show optimal bounds on the relaxation time when the maximum degree is constant.
Theorem 2 ([CLV21]).
Suppose there exists a constant $\eta>0$ such that the system is $\eta$-spectrally independent. Then for all constant $\Delta\geq 2$, for any $n$-vertex graph $G=(V,E)$ of maximum degree $\Delta$, the relaxation time of the Glauber dynamics satisfies:
[…]
@@ 1.5 Main Results
[…]
Therefore, the mixing time satisfies $T_{\mathrm{mix}}=O(n^{2}\log{n})$.
If one further assumes a lower bound on the marginal probability for every vertex and every valid spin assignment then Chen, Liu and Vigoda [CLV21] proved an optimal upper bound on the mixing time.
@@
Suppose there exist constants $\eta>0$ and $b>0$ such that the system is $\eta$-spectrally independent and $b$-marginally bounded. Then for all constant $\Delta\geq 2$, for any $n$-vertex graph $G=(V,E)$ of maximum degree $\Delta$, the mixing time of the Glauber dynamics satisfies:
[…]
@@
In particular, there exists a constant $C(\eta,\Delta,b)$ where the mixing time satisfies
[…]
@@
This is an optimal bound on the mixing time as Hayes and Sinclair [HS07] proved that for any graph with constant maximum degree $\Delta$, the mixing time is $\geq C(\Delta)n\log{n}$.
1.6 Outline of Paper
In the following section (Section 2) we will prove the basic properties of the influence matrix $\Psi$, namely that $\Psi$ is positive semidefinite. In Section 3 we will present the basic facts and definitions regarding Markov chains that are central to the upcoming proofs of the main theorems. The local-to-global theorem of Alev and Lau [AL20] known as the Random Walk Theorem is presented in Section 4, and the local walk, which is closely connected to the influence matrix, is introduced in this section. We then prove Theorem 1 of Anari, Liu, and Oveis Gharan [ALO20], using the Random Walk Theorem, in Section 4.5.
@@ 1.6 Outline of Paper
The Random Walk Theorem is proved in Section 5; here we introduce the up and down processes and explore their connection to the local walk. Then in Section 6 we present the proof of Theorem 2 of Chen, Liu, and Vigoda [CLV21], which shows that spectral independence implies an optimal bound on the spectral gap of the Glauber dynamics. Finally, in Section 7 we present the high-level ideas used to further extend the proof from variance contraction to entropy contraction and obtain optimal mixing time.
@@
The results for generating a random basis of a matroid are presented beginning in Section 8 where we review the relevant definitions for matroids and present the main result for the mixing time of the bases exchange walk. The analysis of the bases exchange walk utilizes the Random Walk Theorem, presented earlier in Section 5, and the Trickle Down Theorem, which is presented in Section 9. We then prove fast mixing of the bases exchange walk in Section 10, and finally we prove the Trickle Down Theorem in Section 11.

### v4

@@ 1.4 Key Definitions: Spectral Independence and the Influence Matrix
For the special case of the hard-core model, the configurations of the model are independent sets. Hence, we denote a configuration by the set of vertices in the independent set $I\subset V$. We refer to those vertices $v\in I$ as occupied and those $w\notin I$ as unoccupied. In this case of the hard-core model by referring to the subset of occupied vertices $I\subset V$ this simplifies the notation. For example, returning to the definition of the influence matrix in Definition 1.1, for the hard-core model on $G=(V,E)$, for $u,v\in V$, the $(u,v)$-entry of the influence matrix $\Psi$ can be compactly expressed as the following:
$\Psi(u\rightarrow v):=\mu(v|u)-\mu(v|\overline{u}),$
where $u,v$ denote the events $u\in I,v\in I$ respectively, and $\overline{u}$ denotes the event $u\notin I$.
As discussed earlier, for the hard-core model, since the configurations correspond to independent sets then a pinning $\tau$ on $S$ can be realized by considering the appropriate induced subgraph $H$ of $G$. Consequently, for a graph $G=(V,E)$ if we show for all induced subgraphs $H$ of $G$ that $\lambda_{\max}(\Psi_{H})\leq 1+\eta$ (without any pinnings) then that implies the spectral independence condition Definition 1.2 holds for $G$ (since every pinning $\tau$ can be realized by an appropriate induced subgraph $H$).
In Section 2 we prove several fundamental properties of the influence matrix $\Psi$, including the fact that the eigenvalues of $\Psi$ are real and non-negative; we also present several basic techniques for bounding the maximum eigenvalue of the influence matrix $\Psi$ in Sections 2.2, 2.3 and 2.4, and introduce the notion of weak spectral independence in Section 2.5.
1.5 Main Results: Fast Mixing via Spectral Independence
Anari, Liu, and Oveis Gharan [ALO20] proved the following result that spectral independence implies polynomial mixing time of the Glauber dynamics.
Theorem 1.4 ([ALO20]).
Suppose $\mu$ is a distribution fully supported on $\Omega\subset\{0,1\}^{n}$ such that the Glauber dynamics for $\mu$ is ergodic. If $\mu$ is $\eta$-spectrally independent, then the relaxation time of the Glauber dynamics for $\mu$ satisfies:
$T_{\mathrm{relax}}\leq O(n^{1+\eta}).$
Therefore, the mixing time satisfies $T_{\mathrm{mix}}=O(n^{1+\eta}\log(1/\mu^{*}))$ where $\mu^{*}:=\min_{\sigma\in\Omega}\mu(\sigma)$.
For the hard-core model with fugacity $\lambda$, note that $\log(1/\mu^{*})=O_{\lambda}(n)$ and hence Theorem 1.4 gives a mixing time upper bound of $O(n^{2+\eta})$.
The above result was improved by Chen, Liu, and Vigoda [CLV21a] to show optimal bounds on the relaxation time when the maximum degree is constant.
Theorem 1.5 ([CLV21a]).
For all constant $\Delta\geq 2$, for any $n$-vertex graph $G=(V,E)$ of maximum degree $\Delta$ and fugacity $\lambda>0$, if the hard-core model on $G$ is $\eta$-spectrally independent for a constant $\eta>0$, then the relaxation time of the Glauber dynamics satisfies:
$T_{\mathrm{relax}}\leq O(n).$
@@ 1.5 Main Results: Fast Mixing via Spectral Independence
$T_{\mathrm{relax}}\leq C(\eta,\Delta)n.$
Therefore, the mixing time satisfies $T_{\mathrm{mix}}=O(n^{2})$.
The above result shows that spectral independence implies optimal relaxation time. In fact, spectral independence is a necessary condition for optimal relaxation time: Anari, Jain, Koehler, Pham, and Vuong [AJK+24] proved that optimal relaxation time of the Glauber dynamics implies spectral independence.
Theorem 1.6.
Let $\mu$ be a distribution on $\{0,1\}^{n}$. If the Glauber dynamics for $\mu$ has relaxation time bounded by $\leq Cn$ then $\lambda_{\max}(\Psi_{\mu})\leq C$. Furthermore, if for any pinning $\tau$ on a subset $S$ of vertices, the Glauber dynamics for $\mu_{\tau}$ has relaxation time bounded by $\leq C\cdot(n-|S|)$, then $\mu$ is (C1)(C-1)-spectrally independent.
If one further assumes a lower bound on the marginal probability for every vertex and every valid pinning, then Chen, Liu and Vigoda [CLV21a] proved an optimal upper bound on the mixing time, rather than the relaxation time.
@@
For all constant $\Delta\geq 2$, for any $n$-vertex graph $G=(V,E)$ of maximum degree $\Delta$ and fugacity $\lambda>0$, if the hard-core model on $G$ is $\eta$-spectrally independent and $b$-marginally bounded for constants $\eta,b>0$, then the mixing time of the Glauber dynamics satisfies:
$T_{\mathrm{mix}}\leq O(n\log{n}).$
@@
In particular, there exists a constant $C(\eta,\Delta,b)$ where the mixing time satisfies for all $\varepsilon>0$,
$T_{\mathrm{mix}}(\varepsilon)\leq C(\eta,\Delta,b)n\log(n/\varepsilon).$
As noted in the theorem statement, Theorem 1.4 holds for any distribution on $\Omega\subseteq\{0,1\}^{n}$ where the Glauber dynamics is ergodic. In contrast, Theorems 1.5 and 1.8 hold for any 2-spin system (i.e., binary graphical model with pairwise interactions), the generalizations to $q>2$ spins are presented in Section 12.2. Examples of 2-spin systems are the hardcore model, the Ising model, and matchings (when viewed as independent sets of the line graph).
@@
As pointed out earlier, Theorem 1.8 establishes an optimal bound on the mixing time of $O(n\log{n})$ as Hayes and Sinclair [HS07] proved that for any graph with constant maximum degree $\Delta$, the mixing time is $\geq C^{\prime}(\Delta)n\log{n}$ for some constant $C^{\prime}(\Delta)$.
Closely connected to the influence matrix is a notion of a local walk, which is introduced in Section 4. The local walk is a Markov chain which is defined on spin assignments for individual vertices in such a way that the spectral gap of the local walk is closely related to the maximum eigenvalue of the influence matrix, and hence the spectral independence property yields a lower bound on the spectral gap of the local walk.
The convergence results stated above (Theorems 1.4, 1.5 and 1.8) are established using a so-called local-to-global theorem of Alev and Lau [AL20], which is known as the Random Walk Theorem, presented in Section 4. The Random Walk Theorem relates the spectral gap of the local walk to the spectral gap of the Glauber dynamics, which corresponds to the global walk in the framework presented in Section 4. Applying the Random Walk Theorem, in Section 4.5, we prove Theorem 1.4 of Anari, Liu, and Oveis Gharan [ALO20]which shows that spectral independence implies polynomial mixing of the Glauber dynamics where the exponent in the polynomial depends on the spectral independence constant.
@@
The Random Walk Theorem is proved in Section 5, where we introduce the associated up and down processes and explore their connection to the local walk. Subsequently, in Section 6, we present the proof of Theorem 1.5 of Chen, Liu, and Vigoda [CLV21a], which shows that spectral independence implies an optimal bound on the spectral gap of the Glauber dynamics; this proof utilizes an improved version of the Random Walk Theorem which is presented in Section 6.2 and proved in Section 6.6. Finally, in Section 6.8 we present the high-level ideas used to further extend the proof from variance contraction to entropy contraction and obtain optimal mixing time. Theorem 1.6 is proved in Section 9.5.
1.6 Application: Hard-core Model
An exciting consequence of the optimal mixing result stated in Theorem 1.8 is a beautiful computational phase transition for the hard-core model. In particular, for every $\Delta\geq 3$ there is a critical point $\lambda_{c}(\Delta)$ where a computational phase transition occurs on graphs of maximum degree $\Delta$. Namely, for all $\lambda<\lambda_{c}(\Delta)$, the Glauber dynamics has optimal mixing time on any graph of maximum degree $\Delta$; this is established using Theorem 1.8. On the other side, when $\lambda>\lambda_{c}(\Delta)$ then it is computationally hard to generate samples from a distribution which is close to the Gibbs distribution on all graphs of maximum degree $\Delta$.
The critical point $\lambda_{c}(\Delta)$ is a consequence of phase transition for the Gibbs distribution on $\Delta$-regular trees. The critical point $\lambda_{c}(\Delta)$ is defined as follows:
$\lambda_{c}(\Delta):=\frac{(\Delta-1)^{\Delta-1}}{(\Delta-2)^{\Delta}}.$
Note, $\lambda_{c}(\Delta)$ is a decreasing function, and for all $\Delta\leq 5$ then $\lambda_{c}(\Delta)>1$, while for $\Delta\geq 6$ then $\lambda_{c}(\Delta)<1$ ($\lambda_{c}(\Delta)=1$ at $\Delta\approx 5.14104$).
The function $\lambda_{c}(\Delta)$ corresponds to the critical point for the following decay of correlations property on trees. Fix an integer $\Delta\geq 3$. For an integer $\ell\geq 1$, let $T_{\ell}$ denote the complete $\Delta$-regular tree of height $\ell$. More precisely, $T_{\ell}$ is the tree where the root is at level 0, all leaves are at level $\ell$, and every non-leaf vertex has $\Delta-1$ children. (Whether the root vertex has degree $\Delta$ or $\Delta-1$ makes negligible difference to the following discussion.)
Let $\mu_{\ell}$ denote the Gibbs distribution on the tree $T_{\ell}$, and let $p_{\ell}$ denote the probability that the root vertex is in an independent set drawn from the distribution $\mu_{\ell}$. Note, the maximum sized independent set in $T_{\ell}$ contains all of the leaves and then every other level (as we move up the tree). Thus, for even $\ell$ the root is in the maximum sized independent set, and for odd $\ell$ the root is not in the maximum sized independent set. Consequently, for large $\lambda$, we expect that $p_{2\ell}>p_{2\ell+1}$ as the root has a bias to be in the independent set for even heights and a bias to be out of the independent set for odd heights.
There is a phase transition on complete $\Delta$-regular trees at $\lambda_{c}(\Delta)$ in the following sense. More formally, for all $\Delta\geq 3$ the following holds:
$\mbox{For all }\lambda\leq\lambda_{c}(\Delta),$ $\lim_{\ell\rightarrow\infty}p_{2\ell+1}=\lim_{\ell\rightarrow\infty}p_{2\ell}$
$\mbox{For all }\lambda>\lambda_{c}(\Delta),$ $\lim_{\ell\rightarrow\infty}p_{2\ell+1}<\lim_{\ell\rightarrow\infty}p_{2\ell}$
This phase transition is derived and explained in more detail in Section 9.1.
In statistical physics terminology, $\lambda_{c}(\Delta)$ is the critical point for the uniqueness/non-uniqueness phase transition on the infinite $\Delta$-regular tree. A Gibbs measure for the hard-core model on the infinite $\Delta$-regular tree is any measure such that for every finite subset of vertices and any pinning outside the subset, the conditional distribution gives a hard-core model with the corresponding pinning. In particular, we can define two extremal Gibbs measures, $\mu_{\mathrm{even}}$ and $\mu_{\mathrm{odd}}$, which can be obtained by taking the limit of finite-volume Gibbs distributions on $T_{\ell}$ with even and odd boundary conditions, respectively, which are fixed assignments (i.e., pinnings) to all of the leaves.
When $\lambda\leq\lambda_{c}(\Delta)$ then the extremal measures are the same $\mu_{\mathrm{even}}=\mu_{\mathrm{odd}}$ and hence there is a unique infinite-volume Gibbs measure; this is called the tree uniqueness region. On the other side, when $\lambda>\lambda_{c}(\Delta)$ then $\mu_{\mathrm{even}}\neq\mu_{\mathrm{odd}}$ and hence there are multiple infinite-volume Gibbs measures; this is called the tree non-uniqueness region. We refer the interested reader to [Geo11, FV17] for a more comprehensive introduction to infinite-volume Gibbs measures and related phase transitions.
The amazing aspect is that the decay of correlations property on $\Delta$-regular trees directly relates to the computational phase transition on all graphs of maximum degree $\Delta$. We can now formally state the optimal mixing result in the tree uniqueness region.
Theorem 1.9.
For all $\Delta\geq 3$, all $\delta>0$, there exists $C(\Delta,\delta)$, for all $\lambda\leq(1-\delta)\lambda_{c}(\Delta)$, for any $n$-vertex graph $G=(V,E)$ of maximum degree $\Delta$, the mixing time of the Glauber dynamics for the hard-core model satisfies for all $\varepsilon>0$,
$T_{\mathrm{mix}}(\varepsilon)\leq C(\Delta,\delta)n\log(n/\epsilon).$
Remark 1.10.
The independent works of Chen, Feng, Yin, and Zhang [CFYZ22] and Chen and Eldan [CE22] improved Theorem 1.9 to obtain $T_{\mathrm{mix}}\leq C(\delta)n\log(n/\varepsilon)$ and thus extend the theorem to classes of graphs where the maximum degree $\Delta$ grows as a function of $n$. We will show in Section 10 (see Remark 10.2) how the tools presented for the proof of the following theorem (Theorem 1.11) yield $T_{\mathrm{relax}}\leq C(\delta)n$, thus extending Theorem 1.9 to unbounded degrees for the relaxation time.
A remarkable result of Chen, Chen, Yin, and Zhang [CCYZ25] proved polynomial mixing time for the Glauber dynamics all the way up to and including the critical point.
Theorem 1.11 ([T]heorem 1.1).
CCYZ-critical-hard-core] There exist constants $C,C^{\prime}>0$, such that for all $\Delta\geq 3$, for any graph $G=(V,E)$ of maximum degree $\Delta$, the relaxation time of the Glauber dynamics for the hard-core model at fugacity $\lambda\leq\lambda_{c}(\Delta)$ satisfies
$T_{\mathrm{relax}}\leq C^{\prime}n^{C}.$
Theorem 1.11 is proved in Section 10.
On the other side, it is computationally hard to approximately sample from the Gibbs distribution and approximate the partition function in the tree non-uniqueness region; in fact, not only is it hard to obtain an $\mathsf{FPRAS}$ (i.e., an arbitrarily close approximation) for the partition function, it is hard to approximate the partition function within some exponential factor even when restricted to triangle-free graphs. The below result is not proved in this monograph, we refer the reader to [GŠV15, CGG+16] for more results in this context (including so-called #BIS-hardness on bipartite graphs in the tree non-uniqueness region).
Theorem 1.12 ([Sly10, SS14, GŠV16, GŠV15]).
For all $\Delta\geq 3$, all $\delta>0$, there exists $\varepsilon=\varepsilon(\Delta,\delta)$, for all $\lambda\geq(1+\delta)\lambda_{c}(\Delta)$, assuming $\mathsf{RP}\neq\mathsf{NP}$, there is no algorithm which for any triangle-free graph $G=(V,E)$ of maximum degree $\Delta$ produces an estimate $\mathsf{EST}$ where:
$\mathrm{Pr}\left(Z_{G,\lambda}2^{-\varepsilon n}\leq\mathsf{EST}\leq Z_{G,\lambda}2^{\varepsilon n}\right)\geq 3/4,$
and runs in time polynomial in $n=|V|$.
The above hardness result is for general graphs, or triangle-free graphs. If we restrict attention to a restricted class of graphs then the threshold for rapid mixing of the Glauber dynamics might occur at a different point. A natural class of graphs to consider are random $\Delta$-regular graphs. Since these graphs are locally tree-like (in particular, they have few short cycles with high probability over the choice of graphs) then we might expect that the tree uniqueness threshold $\lambda_{c}(\Delta)$ is still the threshold for rapid mixing. Surprisingly, Chen, Chen, Chen, Yin and Zhang [CCC+25] showed that when restricted to random $\Delta$-regular graphs then we can obtain fast mixing beyond the uniqueness threshold.
Theorem 1.13 ([T]heorem 1.1).
CCCYZ-random-regular] For all $\Delta\geq 3$, there exists $C(\Delta)$ where the following holds. For $G$ chosen uniformly at random from all $\Delta$-regular graphs, then with probability 1o(1)1-o(1) over the choice of $G$, for all $\lambda<\frac{1}{3\sqrt{\Delta}}$ the Glauber dynamics for the hard-core model on $G$ at fugacity $\lambda$ satisfies:
$T_{\mathrm{mix}}(\varepsilon)\leq C(\Delta)n\log(n/\varepsilon).$
Theorem 1.13 is proved in Section 11. Inspired by Theorem 1.13 it is natural to conjecture that the rapid mixing threshold on random $\Delta$-regular graphs is closely related to the reconstruction threshold, which considers the influence of typical boundary conditions on complete $\Delta$-regular trees, as opposed to worst-case boundaries as considered in the uniqueness threshold; we refer the interested reader to [BST16, RŠV+14] for further discussion regarding the reconstruction threshold in the hard-core model.
We outline results for other spin systems in Section 12.2, including the Ising model results in Section 12.1 and colorings results in Section 12.2.
1.7 Methods for Establishing Spectral Independence
In Section 9 we present general techniques for establishing spectral independence, focusing on three approaches as outlined below. As a consequence, for the hard-core model, we prove that spectral independence holds up to the critical point $\lambda_{c}(\Delta)$ on any graph of maximum degree $\Delta$.
Theorem 1.14.
For all $\Delta\geq 3$, all $\delta>0$, for all $\lambda\leq(1-\delta)\lambda_{c}(\Delta)$, for any $n$-vertex graph $G=(V,E)$ of maximum degree $\Delta$, the Gibbs distribution $\mu_{G,\lambda}$ is $\eta(\delta)$-spectrally independent where $\eta(\delta)\leq 16\sqrt{2}/\delta$.
Combining Theorem 1.14 with the results from Section 1.5, which show that spectral independence implies fast mixing, yields the majority of the fast mixing results for the hard-core model stated in Section 1.6. Towards proving Theorem 1.14, we present several techniques for establishing spectral independence.
In Section 9.2 we present the correlation-decay approach of Weitz [Wei06] and how it yields spectral independence. At a high-level, we show that to upper bound the row-sum of the influence matrix, which is the sum of the influences of a fixed vertex, one can consider influences in a tree, where the tree corresponds to self-avoiding walks in the graph. Then one can use the potential function method to prove that the sum of the influences in the tree decay exponentially fast (in the distance) when the fugacity is in the tree-uniqueness region. In Section 9.3 we show that stability of the partition function, so-called zero-freeness, also implies spectral independence; such conditions were used in the approximate counting algorithm introduced by Barvinok [Bar16a]. Finally, in Section 9.5, we show that a contractive coupling, and more generally optimal relaxation time for a local chain, implies spectral independence.
1.8 Matroids and Trickle-Down Theorem
We utilize the spectral independence approach to establish fast mixing of a Markov chain, known as the bases-exchange walk, for generating a random basis of a matroid, see Section 7 for a more comprehensive introduction and the associated proofs. Anari, Liu, Oveis Gharan, and Vinzant [ALOV19] presented the first proof of fast mixing of the bases-exchange walk for arbitrary matroids. In fact, the work of [ALOV19] for matroids inspired the idea of spectral independence, however we present these results in reverse chronological order. We do not directly utilize spectral independence to prove fast mixing of the bases-exchange walk, instead we utilize the framework of down-up chains and the associated Random Walk Theorem (Theorem 4.1) which are presented in Section 5. One additional tool utilized for the analysis of the bases-exchange walk is the Trickle-Down Theorem of Oppenheim [Opp18] which is used to bound the spectral gap of the local walks considered in the Random Walk Theorem, see Section 8.
We refer the reader to Section 7 for an introduction to matroids and associated terminology, including bases of a matroid. A matroid $\mathcal{M}=(E,\mathcal{I})$ consists of a ground set $E$ and a collection of independent sets $\mathcal{I}$. For the example of a graphic matroid the ground set is the set of edges $E$ of a graph $G=(V,E)$. The independent sets are subsets of the ground set with the requirement that every subset of an independent set is also an independent set (hence the collection $\mathcal{I}$ is downward closed), and the collection of independent sets also need to satisfy the so-called exchange property, see Section 7. The exchange property implies that all maximal independent sets are of the same size; these maximal independents are referred to as the bases of the matroid and their size is the rank of the matroid. For the example of a graphic matroid, the independent sets are acyclic subgraphs (viewed as subsets of edges) of a graph $G=(V,E)$ and the bases are spanning forests of $G$.
The bases-exchange walk is a Markov chain on the bases of a matroid where the transitions utilize the exchange property in the following manner: from a basis $B_{t}$ at time $t$, we choose an element $e\in B_{t}$ uniformly at random, and an element $f$ uniformly from those elements such that $B^{\prime}=B_{t}\cup\{f\}\setminus\{e\}$ is a basis, and then we set $B_{t+1}=B^{\prime}$ as the state at time $t+1$. The main result presented in Section 7 is a bound on the relaxation time of the bases-exchange walk.
Theorem 1.15 ([ALOV19]).
For a matroid $\mathcal{M}=(E,\mathcal{I})$, the relaxation time of the bases-exchange walk satisfies $T_{\mathrm{relax}}\leq r(\mathcal{M})$, where $r(\mathcal{M})$ is the rank of the matroid. Consequently, the mixing time of the bases-exchange walk satisfies $T_{\mathrm{mix}}=O(r(\mathcal{M})\log(|\mathcal{I}|))=O(r(\mathcal{M})^{2}\log n)$ where $n=|E|$.
The mixing time bound was improved by Cryan, Guo, and Mousa [CGM21] to $O(r(\mathcal{M})\log{\log(|\mathcal{I}|)})=O(r(\mathcal{M})\log r(\mathcal{M})+r(\mathcal{M})\log{\log{n}})$; moreover, the $\log\log n$ term was subsequently removed in [ALO+21], achieving $O(r(\mathcal{M})\log r(\mathcal{M}))$ mixing time.
@@ 1.8 Matroids and Trickle-Down Theorem
The results for generating a random basis of a matroid are presented beginning in Section 7 where we review the relevant definitions for matroids and present the main result for the mixing time of the bases-exchange walk. The analysis of the bases-exchange walk is presented in Section 8; the proof utilizes the Random Walk Theorem, presented earlier in Section 5, and the Trickle-down Theorem, which is presented in Section 8.2. The Trickle-down Theorem is used to obtain an optimal bound on the spectral gap of the associated local walks. The proof of fast mixing of the bases-exchange walk is presented in Section 8.3, and the proof of the Trickle-down Theorem is presented in Section 8.5.